OpenStudy (anonymous):

A 60 lb girl and a 50 lb boy face each other on frictionless rollerskates. The girl pushes the boy, who moves away at a speed of 4.0 ft/s. The girl's speed is?

OpenStudy (anonymous):

M1 * V1 = M2* V2 60lb(v)=50lb(-4.0ft/s).....- velocity man in opposite direction v is the velocity of the girl =(10/3)ft/s